3
The outbreak of Bird Flu in 1997 resulted in the Hong Kong government ordering the culling of more than 1.5 million chickens. The culling of chickens was simultaneously accompanied by consumers reducing their demand for life chickens due to the bird flu. Using demand and supply analysis, what was the impact on price and quantity in the market for life chickens?

4
Assume the price of a good increase from $6 to $8, leading to a fall in quantity demanded from 50 to 40 units. Calculate the price elasticity of demand for the good at this price range and explain how total revenue will be impacted by the increase in price?
1
Expert's answer
2016-05-06T12:24:02-0400
3. If the culling of chickens was simultaneously accompanied by consumers reducing their demand for life chickens due to the bird flu, then according to decrease in supply and demand (both curves shift leftward) equilibrium price decreases, increases or may not change and equilibrium quantity decreases in the market for chickens.
4. If the price of a good increase from $6 to $8, leading to a fall in quantity demanded from 50 to 40 units, then the price elasticity of demand is: Ed = (40 - 50)/(8 - 6)*(8 + 6)/(40 + 50) = -10/2*14/90 = -7/9, |Ed| < 1, so the demand is inelastic. It means that total revenue will decrease according to the increase in price.